Question 13 Snow Cap Springs produces and sells water filtration systems for homeowners. Information regarding its three models is shown below. Basic Basic Plus Premium Total Units sold 720 240 240 1,200 Selling price $251 $405 $804 Variable cost $190 $291 $417 The company’s total fixed costs to produce the filtration systems are $205,200. Determine the sales mix as a function of units sold for the three products. Basic Basic Plus Premium The sales mix as a function of units sold % % % LINK TO TEXT Determine the weighted-average unit contribution margin. (Round answer to 2 decimal places, e.g. 10.25.) Weighted-average unit contribution margin $ LINK TO TEXT Determine the total number of units that the company must produce to break even. Break-even point in units units LINK TO TEXT Determine the number of units of each model that the company must produce to break even. Basic units Basic Plus units Premium units Click if you would like to Show Work for this question: Open Show Work
Answer #1
1
Sales mix as a function of units sold:
Basic 60% =720/1200
Basic Plus 20% =240/1200
Premium 20% =240/1200
2
Basic Basic Plus Premium Total
Selling price 251 405 804
Less: Variable cost 190 291 417
Unit Contribution margin 61 114 387
X Sales mix percentage 60% 20% 20%
Weighted Unit Contribution margin 36.60 22.80 77.40 136.80
Weighted-average unit contribution margin 136.80
3
Break-even point in units 1500 units =205200/136.8
4
Basic 900 units =1500*60%
Basic Plus 300 units =1500*20%
Premium 300 units =1500*20%